# How to Beat Nakamoto in the Race

### v1

@@ Abstract.
This paper studies proof-of-work Nakamoto consensus under bounded network delays, settling two long-standing questions in blockchain security: How can an adversary most effectively attack block safety under a given block confirmation latency? And what is the resulting probability of safety violation? A Markov decision process (MDP) framework is introduced to precise characterize the system state (including the tree and timings of all blocks mined), the adversary’s potential actions, and the state transitions due to the adversarial action and the random block arrival processes. An optimal attack, called bait-and-switch, is proposed and proved to maximize the adversary’s chance of violating block safety by “beating Nakamoto in the race”. The exact probability of this violation is calculated for any confirmation depth using Markov chain analysis, offering fresh insights into the interplay of network delay, confirmation rules, and blockchain security.
@@
doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; October 13–17, 2025; Taipei, Taiwanisbn: ?$c$s: Security and privacy Distributed systems security
@@ 1. Introduction
The Nakamoto consensus, underpinned by the longest-chain fork-choice rule, serves as a foundational pillar in blockchain technology, facilitating trustless and decentralized consensus within Bitcoin and numerous other systems. This pioneering mechanism has demonstrated consistency and liveness both theoretically and in practice, enabling consensus without a central authority. Provided that a sufficient majority of the network adheres to the consensus rules, the Nakamoto consensus ensures that the probability of a block’s safety (or blockchain consistency) being violated vanishes as it gains more confirmations within a longest chain.
While the safety of Nakamoto consensus is well understood when all honest nodes are always fully synchronized with identical views (nakamoto2008bitcoin, ; garay2015bitcoin, ), its precise safety under weaker synchrony assumptions and finite confirmation latency, crucial for real-world scenarios, remains an open question. Garay et al. (garay2015bitcoin, ) conducted a formal security analysis of Bitcoin’s backbone protocol, followed by diverse approaches in later works (bagaria2019prism, ; pass2017analysis, ; kiffer2018better, ; gazi2020tight, ; garay2020full, ; li2021close, ; gazi2022practical, ; gazi2023practical, ; cao2025security, ; doger2024refined, ), including renewal process races (li2021close, ; cao2025security, ), Markov chains (kiffer2018better, ; doger2024refined, ), and numerical evaluation using dynamic programming (gazi2022practical, ; gazi2023practical, ). However, previous works only offer upper and lower bounds on block safety, where the gap between those bounds can be very large when the adversarial power is near the ultimate fault tolerance. Notably, no attack has been identified as optimal under network delays, leaving a critical aspect of the problem unresolved.
This paper adopts the widely adopted $\Delta$-synchrony model, which assumes that any block or message generated or received by an honest node is delivered to all other honest nodes within a fixed delay of $\Delta$ seconds. In the proof-of-work (PoW) (nakamoto2008bitcoin, ) setting, we propose an attack, termed “bait-and-switch”, and prove its optimality under the $k$-confirmation commitment rule. In this attack, the adversary privately maintains a fork and delays the propagation of honest blocks for the maximum allowed duration ($\Delta$). At a strategic moment—specifically, when the private chain is behind but matches the length of the public chain visible to all honest nodes—the adversary reveals the private fork as “bait”. This tactic causes honest nodes to perceive the adversary’s chain as a longest and redirect their efforts to extend it. In effect, the adversary opportunistically harnesses honest mining power to increase their chances of winning Nakamoto’s longest-chain race. Although the adversary has countless possible strategies, we prove that no other attack offers a better chance of “beating Nakamoto in the race”!
@@
It is commonly understood that the adversary’s problem is to determine how to allocate resources and influence honest participants at every moment in time based on all information available until then. While prior work has modeled selfish mining using Markov decision processes (MDPs) (sapirshtein2017optimal, ; bai2023blockchain, ; feng2019selfish, ; zur2020efficient, ; doger2025when, ), the rapid expansion of the state space has been a major roadblock. Suboptimal learning algorithms have been proposed to yield a solution with constrained complexity (bar2022werlman, ). In this paper, we fully develop a general infinite-horizon MDP framework for the Nakamoto consensus and prove for the first time that a Markov definite policy (or attack) maximizes the probability of eventual safety violation. We also reduce the state representation—from a full blocktree and per-node views—to a compact, variable-length tuple that retains all the necessary information for assessing block safety. Using this streamlined framework, the bait-and-switch attack is proved to be an optimal policy.
@@ 2. Model
We first introduce a succinct notation for referring to blocks and chains.
@@
A sequence of $n+1$ block numbers, $(b_{0},b_{1},\dots,b_{n})$, defines a blockchain (or simply a chain) if $b_{0}=0$ and block $b_{i-1}$ is block $b_{i}$’s parent for every $i=1,\dots,n$. Since every chain can be uniquely identified by its last block, we also refer to the preceding chain as chain $b_{n}$. As a convention, we say the height of this chain is $n$. Also, we say block $b_{i+1}$ extends chain $b_{i}$ and that chain $b_{i}$ is a prefix of chain $b_{n}$ for every $i=0,\dots,n-1$. In general, we use $h_{b}$ to denote the height of chain $b$, which is also referred to as the height of block $b$. Evidently, the genesis block is the only block on height 0.
@@
Figure 1. An example of blocks, and chains.
@@ 2.1. Decentralized Blockchain System
A Nakamoto consensus protocol enables pseudonymous participants to achieve probabilistic state replication with high reliability (garay2015bitcoin, ; nakamoto2008bitcoin, ). We assume a well decentralized system with a large number of participants (nodes or miners). Without loss of generality, we consider a single adversarial participant, since multiple adversaries can be treated as a single coordinated entity—yielding a more powerful adversary. The protocol can be characterized by the following key components and principles.
@@ 2.1.1. Proof of Work
Nodes compete to solve cryptographic puzzles, and each successful solution grants the right to propose a new block extending an existing chain. The puzzle—essentially a brute-force search for a solution to a hash inequality based on SHA-256—is memoryless nature. Since nodes typically perform an enormous number of independent attempts per second, the block arrival process is well modeled as a continuous-time Poisson point process (see, e.g., (guo2022bitcoin, ; dembo2020everything, )).
@@ 2.1.2. Longest-Chain Fork Choice
When an honest node gets the opportunity to propose a block, the node always extends a highest valid chain in its view. In the event that multiple highest chains exist in the view, we adopt a worst-case assumption, allowing the adversary to resolve the tie in any manner it prefers for that node.
@@ 2.1.3. Block Propagation Delay
When an honest node mines a new block or receives a new valid block for the first time, it relays the block to its peers through a peer-to-peer network. Due to network delays, different honest nodes may maintain different views of the blockchain. The widely used $\Delta$-synchrony model posits an upper bound of $\Delta$ seconds on end-to-end communication delays among nodes, such that any information known to an honest node at time $t$ becomes known to all honest nodes by time $t+\Delta$.
@@ 2.1.4. Commitment rule
This paper adopts the $k$-confirmation commitment rule, namely, a node commits to a block when at least k1k-1 blocks have been mined on top of it as part of a longest chain known to that node, where $k$ is an integer selected by the node. If, after a block at height $\eta$ is committed by some honest node, it remains permanently included in all longest chains in all honest nodes’ views, then the block—and thus height $\eta$—is (forever) safe; otherwise the safety of the block (and its height) is violated (at some point).
@@ 2.1.5. Adversarial Capabilities
  • It may withhold its own chains for an arbitrary duration and selectively reveal any chain to any honest node at any time, within the synchrony bounds.
@@ 2.2. Enhanced Adversary
The enhancement appears subtle: For a block that arrives at $t$, the original adversary must decide which node(s) and chain(s) are the potential miner(s) and parent(s) by allocating (adversarial and honest) mining power based on available information in $[0,t)$. In contrast, the enhanced adversary may choose both the miner and the parent of the block upon its arrival, based on all available information in $[0,t]$ (which is more than $[0,t)$). In particular, if the original adversary splits its mining power between two chains, it cannot predict which one will be extended, whereas the enhanced adversary can decide which one is extended upon the block’s arrival.
@@ 2.3. A Tree-based Mathematical Model
Following Definitions 1 and 2, we describe a succinct mathematical model for Nakamoto’s blockchain system with an enhanced adversary, termed tree-based model.
@@
A block or chain may become public at a specific point in time strictly after its creation and will remain public at all times afterward. A chain being public at $t$ is equivalent to all of its blocks being public at $t$. As a convention, the genesis block becomes public at time 0 and block 1 is mined afterward, i.e., $t_{1}>0$. At time $t$, the height of a highest public chain is referred to as the public height.
@@
In Fig. 1, at $t_{10}$, nonpublic blocks (7, 9, and 10) are shaded. The public height is 3. Chains 5, 7, 8, 9, and 10 are $t_{10}$-credible. No other chain in Fig. 1 is $t_{10}$-credible.
A chain known to all honest nodes is termed public, per Definition 3. A chain that is at least as high as any public chain is deemed credible, according to Definition 4. At any given time, every existing chain comprises a public prefix, possibly extended by nonpublic blocks.
@@
By convention, the genesis block is neither an A-block or an H-block. Each subsequent block is either an A-block or an H-block. An H-block that arrives at $t$ must be strictly higher than the public height at time $t$. For future convenience, we let $\bm{A}_{t}$ (resp. $\bm{H}_{t}$) denote the index set of all A-blocks (resp. H-blocks) that arrive during $(0,t]$.
In the example in Fig. 1, each A-block is marked by a pair of horns on top: $\{2,7\}\subset\bm{A}_{t_{10}}$ and $\{1,3,4,5,6,8,9,10\}\subset\bm{H}_{t_{10}}$.
A-blocks and H-blocks represent blocks mined by honest and adversarial nodes, respectively. The adherence of honest nodes to the longest-chain fork-choice rule is encapsulated by the requirement that each new H-block must be higher than all existing public chains, as specified in Definition 5. In contrast, an A-block can have any existing block as its parent. Throughout this paper, the terms “longest” and “highest” are used interchangeably.
@@
If there exists a $t$-credible chain $c$ for some $t>0$, which includes a block $b$, where $h_{c}\geq h_{b}+k-1$, then we say chain $c$ confirms block $b$ by depth $k$, or block $b$ is $k$-confirmed by chain $c$.
For example, in Fig. 1, chain 7 confirms block 1 by depth 4.
@@
If there ever exist times $t>0$ and $t^{\prime}\geq t>0$, when one $t$-credible chain confirms block $c$ and another $t^{\prime}$-credible chain confirms a different block on the same height, $h_{c}$, we say the safety of height $h_{c}$ is violated, or a safety violation occurs at height $h_{c}$ at time $t^{\prime}$.
Evidently, a longest chain in a node’s view must be credible (Definition 7). For example, in Fig. 1, with $k=3$, block 2 is committed via a $t_{8}$-credible chain, chain 8, as two blocks are mined on top of it within chain 8, which is a longest chain visible to an honest node. However, block 2 is absent from the subsequent $t_{9}$-credible chain (chain 9), where $t_{9}>t_{8}$, resulting in a safety violation at height 1.
@@
Let $H_{t}$ (resp. $A_{t}$) denote the number of honest (resp. adversarial) arrivals during $(0,t]$. Then $(H_{t})_{t>0}$ and $(A_{t})_{t>0}$ are independent Poisson point processes with rates $h$ and $a$, respectively. An honest arrival at $t$ corresponds to the mining of an H-block at $t$. An adversarial arrival corresponds to the mining of an A-block.
@@
2.4. Sufficiency of the Tree-based Model
Evidently, the decentralized system described in Sec. 2.1 is view-based in the sense that its evolution can be fully described using the trajectory of all nodes’ views (including the adversary’s). In the literature, nodes are often treated as interactive Turing machines which follows an algorithm (see, e.g., (garay2015bitcoin, ; garay2020full, )). We next show that the tree-based mathematical model introduced in Sec. 2.3 is sufficient for studying safety violations in the view-based model.
@@ 2.4. Sufficiency of the Tree-based Model
For every realization (or sample path) of the A- and H-block arrival processes, if the enhanced adversary in the view-based model can produce a certain blocktree, yielding a safety violation, so can the adversary in the tree-based mode, and vice versa.
@@
We prove by induction on the sequence of block arrivals. At time $t_{0}=0$, the two models share the same blocktree, which includes only the genesis block. In the view-based model, all honest nodes’ views include only this single-node blocktree. In the tree-based model, the genesis is public.
Suppose the two models share the same blocktree immediately after the arrival time of block (b1)(b-1), and that the public prefix of the blocktree in the tree-based model is included in all honest views in the view-based model. As a consequence, every honest view includes at least one chain that is no lower than the public height.
First, if an honest node mines block $b$ to extend chain $c$ in the view-based model, then chain $c$ is no lower than the public height, so that block $b$ can be made a child of block $c$ in the tree-based model.
Second, by delaying the propagation of all honest blocks maximally in the view-based system, all honest nodes hold the same (public) view except for those who have mined blocks within the most recent $\Delta$ period, which are yet to become public. As a consequence, if an H-block $b$ arrives to extend chain $c$ in the tree-based model, chain $c$ must be credible (with respect to the public view), which implies that chain $c$ is as high as a highest chain in some honest views in the view-based model, so that the enhanced adversary can pick such an honest node as the miner of block $b$ to extend chain $c$.
The preceding arguments guarantee that the view- and tree-based models share the same blocktree immediately after the arrival of block $b$. Moreover, the public prefix of the blocktree in the tree-based model is included in all honest views in the view-based model.
Consider a safety violation in the tree-based model, where one $t$-credible chain $c$ confirms block $b$ on height $s$ and another $t^{\prime}$-credible chain $c^{\prime}$ confirms a different block $b^{\prime}$ on height $s$. Evidently, chain $c$ can be made a highest in some honest view at time $t$ in the view-based model, and $c^{\prime}$ can be made a highest in another honest view at time $t^{\prime}$, creating a safety violation in the view model with the same blocktree. It is easy to see that the converse is also true. Hence the two models are equivalent in the sense that if the enhanced adversary can create a safety violation in the view-based model, then so can the adversary in the tree-based model, and vice versa. ∎

### v2

@@ Abstract.
This paper studies proof-of-work Nakamoto consensus protocols under bounded network delays, settling two long-standing questions in blockchain security: What is the most effective attack on block safety under a given block confirmation latency? And what is the resulting probability of safety violation? A Markov decision process (MDP) framework is introduced to precisely characterize the system state (including the blocktree and timings of all blocks mined), the adversary’s potential actions, and the state transitions due to the adversarial action and the random block arrival processes. An optimal attack, called bait-and-switch, is proposed and proved to maximize the adversary’s chance of violating block safety by “beating Nakamoto in the race”. The exact probability of this violation is calculated for any given confirmation depth using Markov chain analysis, offering fresh insights into the interplay of network delay, confirmation rules, and blockchain security.
@@
$c$s: Security and privacy Distributed systems security
@@ 1. Introduction
The Nakamoto consensus, underpinned by the longest-chain fork-choice rule, serves as a foundational pillar in blockchain technology, facilitating trustless and decentralized consensus in Bitcoin and numerous other systems. This pioneering mechanism has demonstrated consistency and liveness both theoretically and in practice, enabling consensus without a central authority. Provided that a sufficient majority of the network adheres to the consensus protocols, the Nakamoto consensus ensures that the probability of a block’s safety (or blockchain consistency) being violated vanishes as it gains more confirmations within a longest chain.
While the safety of Nakamoto consensus is well understood when all honest nodes are always fully synchronized with identical views (Nakamoto, 2008; Garay et al., 2015), its precise safety under weaker synchrony assumptions and finite confirmation latency, crucial for real-world scenarios, remains an open question. Garay et al. (Garay et al., 2015) conducted a formal security analysis of Bitcoin’s backbone protocol, followed by diverse approaches in later works (Bagaria et al., 2019; Pass et al., 2017; Kiffer et al., 2018; Gaži et al., 2020; Garay et al., 2020; Li et al., 2021; Gaži et al., 2022, 2023; Cao and Guo, 2025; Doger and Ulukus, 2025a), including renewal process races (Li et al., 2021; Cao and Guo, 2025), Markov chains (Kiffer et al., 2018; Doger and Ulukus, 2025a), and numerical evaluation using dynamic programming (Gaži et al., 2022, 2023). However, previous works only offer upper and lower bounds on block safety, where the gap between those bounds can be very large when the adversarial power is near the ultimate fault tolerance. Notably, no attack has been identified as optimal under network delays, leaving a critical aspect of the problem unresolved.
This paper adopts the widely used $\Delta$-synchrony model, which assumes that any block or message generated or received by an honest node is delivered to all other honest nodes within a fixed delay of $\Delta$ seconds. In the proof-of-work (PoW) (Nakamoto, 2008) setting, we propose an attack, termed “bait-and-switch”, and prove its optimality under the $k$-confirmation commitment rule. In this attack, the adversary privately maintains a fork and delays the propagation of honest blocks for the maximum allowed duration ($\Delta$). At a strategic moment—specifically, when the private chain is behind but matches the length of a highest public chain visible to all honest nodes—the adversary reveals the private fork as “bait”. This tactic causes honest nodes to perceive the adversary’s chain as a longest and redirect their efforts to extend it. In effect, the adversary opportunistically harnesses honest mining power to increase their chances of winning Nakamoto’s longest-chain race. Although the adversary has countless possible strategies, we prove that no other attack offers a better chance of “beating Nakamoto in the race”!
@@
It is commonly understood that the adversary’s problem is to determine how to allocate resources and influence honest participants at every moment in time based on all information available until then. While prior work has modeled selfish mining using Markov decision processes (MDPs) (Sapirshtein et al., 2017; Bai et al., 2023; Feng and Niu, 2019; Zur et al., 2020; Doger and Ulukus, 2025b), the rapid expansion of the state space has been a major roadblock. Suboptimal learning algorithms have been proposed to yield a solution with constrained complexity (Bar-Zur et al., 2022). In this paper, we fully develop a general infinite-horizon MDP framework for the Nakamoto consensus and prove for the first time that a Markov definite policy (or attack) maximizes the probability of eventual safety violation. We also reduce the state representation—from a full blocktree and per-node views—to a compact, variable-length tuple that retains all the necessary information for assessing block safety. Using this streamlined framework, the bait-and-switch attack is proved to be an optimal policy.
@@ 2. Model
We introduce a succinct notation for referring to blocks and chains.
@@
A sequence of $n+1$ block numbers, $(b_{0},b_{1},\dots,b_{n})$, defines a blockchain (or simply a chain) if $b_{0}=0$ and block $b_{i-1}$ is block $b_{i}$’s parent for every $i=1,\dots,n$. Since every chain can be uniquely identified by its last block, we also refer to the preceding chain as chain $b_{n}$. As a convention, we say the height of this chain is $n$. Also, we say block $b_{i+1}$ extends chain $b_{i}$ and that chain $b_{i}$ is a prefix of chain $b_{n}$ for every $i=0,\dots,{n}$. In general, we use $h_{b}$ to denote the height of chain $b$, which is also referred to as the height of block $b$. Evidently, the genesis block is the only block at height 0.
@@
Figure 1. An example blocktree.
@@ 2.1. Decentralized Blockchain System
A Nakamoto consensus protocol enables pseudonymous participants to achieve probabilistic state replication with high reliability (Garay et al., 2015; Nakamoto, 2008). We assume a well decentralized system with a large number of participants (nodes or miners). Without loss of generality, we consider a single adversarial participant, since multiple adversaries can be treated as a single coordinated entity—yielding a more powerful adversary. The protocol can be characterized by the following key components and principles.
@@ 2.1.1. Proof of Work
Nodes compete to solve cryptographic puzzles, and each successful solution grants the right to propose a new block extending an existing chain. The puzzle—essentially a brute-force search for a solution to a hash inequality based on SHA-256—is memoryless nature. Since nodes typically perform an astronomical number of independent attempts per second, the block arrival process is well modeled as a continuous-time Poisson point process (see, e.g., (Guo and Ren, 2022; Dembo et al., 2020)).
@@ 2.1.2. Longest-Chain Fork Choice
When an honest node gets the opportunity to propose a block, the node always extends a highest valid chain in its view. In the event that multiple highest chains exist in the view, we adopt a worst-case assumption, allowing the adversary to resolve the tie in any manner it prefers for that node. Throughout this paper, the terms “longest” and “highest” are used interchangeably.
@@ 2.1.3. Block Propagation Delay
When an honest node mines a new block or receives a new valid block for the first time, it relays the block to its peers through a peer-to-peer network. Due to network delays, different honest nodes may maintain different views of the blocktree. The widely used $\Delta$-synchrony model posits an upper bound of $\Delta$ seconds on end-to-end communication delays among nodes, such that any information known to an honest node at time $t$ becomes known to all honest nodes by time $t+\Delta$.
@@ 2.1.4. Commitment rule
This paper adopts the $k$-confirmation commitment rule, namely, a node commits to a block when at least k1k-1 blocks have been mined on top of it as part of a longest chain in that node’s view, where $k$ is an integer selected by the node. If, after a block at height $\eta$ is committed by some honest node, it remains permanently included in all longest chains in all honest nodes’ views, then the block—and thus height $\eta$—is (forever) safe; otherwise the safety of the block (and its height) is violated (at some point).
@@ 2.1.5. Adversarial Capabilities
  • It may withhold its own chains for an arbitrary duration and selectively reveal any chain to any honest node at any time.
@@ 2.2. Enhanced Adversary
The enhancement appears subtle: For a block that arrives at $t$, the original adversary must decide which node(s) and chain(s) are the potential miner(s) and parent(s) by allocating (adversarial and honest) mining power based on available information strictly before $t$. In contrast, the enhanced adversary may choose both the miner and the parent of the block upon its arrival, based on all available information in $[0,t]$ (which is more than $[0,t)$). In particular, if the original adversary splits its mining power between two chains, it cannot predict which one will be extended, whereas the enhanced adversary can decide which one is extended upon the block’s arrival.
@@ 2.3. A Tree-based Mathematical Model
Following Definitions 1 and 2, we introduce a succinct tree-based mathematical model for the view-based blockchain system with an enhanced adversary described in Secs. 2.1 and 2.2.
@@
A block or chain may become public at a specific point in time strictly after its creation and will remain public at all times afterward. Block $b$ is public if and only if chain $b$ is public. A chain being public at $t$ is equivalent to all of its blocks being public at $t$. As a convention, the genesis block becomes public at time 0 and block 1 is mined afterward, i.e., $t_{1}>0$. At time $t$, the height of a highest public chain is referred to as the public height.
At any given time, every existing chain comprises a public prefix, possibly extended by nonpublic blocks. As an example, in Fig. 1, at $t_{10}$, all nonpublic blocks (7, 9, and 10) are shaded. It is easy to see that the public height is 3.
@@
In Fig. 1, chains 5, 7, 8, 9, and 10 are $t_{10}$-credible. No other chain in Fig. 1 is $t_{10}$-credible.
@@
By convention, the genesis block is neither an A-block or an H-block. Each subsequent block is either an A-block or an H-block. An H-block that arrives at $t$ must be strictly higher than the public height at time $t$.
In Fig. 1, each A-block is distinguished by a pair of horns on top. Blocks 2 and 7 are A-blocks; blocks 1, 3–6, and 8–10 are H-blocks.
@@
If there exists a $t$-credible chain $c$ for some $t>0$ that includes a block $b$, where the heights satisfy $h_{c}\geq h_{b}+k-1$, then we say chain $c$ confirms block $b$ at depth $k$, or block $b$ is $k$-confirmed by chain $c$.
In Fig. 1, chain 7 confirms block 1 at depth 4, even though block 7 is nonpublic. At the same time, chain 8 confirms block 2 at depth 3, even though chain 8 is not the longest.
@@
If there ever exist times $t>0$ and $t^{\prime}\geq t$, when one $t$-credible chain confirms block $c$ and another $t^{\prime}$-credible chain confirms a different block at the same height, $h_{c}$, we say the safety of height $h_{c}$ is violated, or a safety violation occurs at height $h_{c}$ at time $t^{\prime}$.
In Fig. 1, only block 1 is 5-confirmed by $t_{10}$, thus there is no violation yet under a 5-confirmation rule. However, if a 4-confirmation rule is adopted instead, then block 7 confirms block 1 at $t_{7}$ and block 10 confirms block 2 at $t_{10}$, causing a violation at height 1.
@@
Let $H_{t}$ (resp. $A_{t}$) denote the number of honest (resp. adversarial) arrivals during $(0,t]$. Let $(H_{t})_{t>0}$ and $(A_{t})_{t>0}$ be independent Poisson point processes with rates $h$ and $a$, respectively. An honest arrival at $t$ corresponds to the mining of an H-block at $t$. An adversarial arrival corresponds to the mining of an A-block.
@@
2.4. Tree-Based Model for View-Based System
The decentralized system described in Secs. 2.1 and 2.2 is view-based in the sense that its evolution can be fully described using the trajectory of all nodes’ views (including the adversary’s). In the literature, nodes are often treated as interactive Turing machines which follows an algorithm (see, e.g., (Garay et al., 2015, 2020)). We next draw connections between the view-based system and its tree-based model.
A chain known to all honest nodes in the view-based system is termed public in the tree-based model, per Definition 3. A chain that is at least as high as any public chain is deemed credible, per Definition 4. A-blocks and H-blocks in the tree-based model represent blocks mined by honest and adversarial nodes, respectively, in a view-based system. The adherence of honest nodes to the longest-chain fork-choice rule is encapsulated by the requirement (in the tree-based model) that each new H-block must be higher than all existing public chains, as specified in Definition 5. In contrast, an A-block can have any existing block as its parent.
Evidently, a longest chain in a node’s view (in a view-based system) is credible in the corresponding tree-based model (Definition 7). For example, in Fig. 1, with $k=3$, block 2 is committed via a $t_{8}$-credible chain, chain 8, as two blocks are mined on top of it within chain 8, which is a longest chain visible to an honest node. However, block 2 is absent from the subsequent $t_{9}$-credible chain (chain 9), where $t_{9}>t_{8}$, resulting in a safety violation at height 1.
We next show that the tree-based mathematical model is sufficient for studying safety violations in the view-based system.
@@ 2.4. Tree-Based Model for View-Based System
For every realization (or sample path) of the A- and H-block arrival processes, if the enhanced adversary in the view-based system can produce a certain blocktree, yielding a safety violation, so can the adversary in the tree-based mode, and vice versa.
@@
We prove by induction on the sequence of block arrivals. At time $t_{0}=0$, the two models share the same blocktree, which includes only the genesis block. In the view-based system, all honest nodes’ views include only this single-node blocktree. In the tree-based model, the genesis is public.
Suppose the two models share the same blocktree immediately after the arrival time of block (b1)(b-1), and that the public prefix of the blocktree in the tree-based model is included in all honest views in the view-based system. As a consequence, every honest view includes at least one chain that is no lower than the public height.
First, if an honest node mines block $b$ to extend chain $c$ in the view-based system, then chain $c$ is no lower than the public height, so that block $b$ can be made a child of block $c$ in the tree-based model.
Second, by delaying the propagation of all honest blocks maximally in the view-based system, all honest nodes hold the same (public) view except for those who have mined blocks within the most recent $\Delta$ period, which are yet to become public. As a consequence, if an H-block $b$ arrives to extend chain $c$ in the tree-based model, chain $c$ must be credible (with respect to the public view), which implies that chain $c$ is as high as a highest chain in some honest views in the view-based system, so that the enhanced adversary can pick such an honest node as the miner of block $b$ to extend chain $c$.
The preceding arguments guarantee that the view-based system and the tree-based model share the same blocktree immediately after the arrival of block $b$. Moreover, the public prefix of the blocktree in the tree-based model is included in all honest views in the view-based system.
Consider a safety violation in the tree-based model, where one $t$-credible chain $c$ confirms block $b$ at height $s$ and another $t^{\prime}$-credible chain $c^{\prime}$ confirms a different block $b^{\prime}$ at height $s$. Evidently, chain $c$ can be made a highest in some honest view at time $t$ in the view-based system, and $c^{\prime}$ can be made a highest in another honest view at time $t^{\prime}$, creating a safety violation in the view model with the same blocktree. It is easy to see that the converse is also true. Hence the two models are equivalent in the sense that if the enhanced adversary can create a safety violation in the view-based system, then so can the adversary in the tree-based model, and vice versa. ∎